The following are two distinct demand functions: 𝑞1=24−0.2𝑃1 and 𝑞2=10−0.05𝑃2. The average cost AC=40+35/q. What price the firm will change with discrimination?


1
Expert's answer
2021-05-27T18:56:54-0400

"Q_1=24-0.2P_1"

"P_1=\frac{24-Q_1}{0.2}"

"Q_2=10-0.05P_2"

"P_2=\frac{10-Q_2}{0.05}"

Average cost; "AC=40+\frac{35}{Q}"

Total Revenue; "TR=P\times Q"

so,

"TR_1=P_1\times Q_1"

"=\frac{24-Q_1}{0.2}\times Q_1"

"=120Q_1-5Q_1^2"

"TR_2=P_2\times Q_2"

"=\frac{10-Q_2}{0.05}\times Q_2"

"=200Q_2-20Q_2^2"

marginal revenue=First derivative of total revenue

"MR_1=\frac{dTR_1}{dQ_1}"

"=120-10Q_1"


"MR_2=\frac{dTR_2}{dQ_2}"

"=200-40Q_2"


Average cost; "AC=(40-\frac{35}{Q})\times Q"

"=40Q-35"

Marginal cost = First derivative of total cost


"MC=\frac{dTC}{dQ}"

"=40"


Profit maximizing quantity of the firm will be when;

"MR_!=MR_2=MC"

when;

"MR_1=MC\\"

"120-10Q_1=40\\10Q_1=120-40\\Q_1=8"


then; "P_1=\frac{24-Q_1}{0.2}"


"=\frac{24-8}{0.2}"


"=80"


and when

"MR_2=MC\\200-40Q_2=40\\40Q_2=200-40\\Q_2=\frac{160}{40}\\Q_2=4"

then

"P_2=\frac{10-Q_1}{0.05}\\"

"=\frac{10-4}{0.05}"


"=120"


Therefore the prices firm will charge with discrimination are "P_1=80\space and\space P_2=120"